Two astronauts are 2.0 m apart in their spaceship. One speaks to the other. The conversation is transmitted to earth via electromagnetic waves. The time it takes for sound waves to travel at 348 m/s through the air between the astronauts equals the time it takes for the electromagnetic waves to travel to the earth. How far away from the earth is the spaceship?   m
Answer #1

This is comparing the speed of sound to the speed of light.

D = R*T for both sound and light

For sound
2.0m = 348 m/s * T
so
T = 2.0/ 348 seconds = 0.005747 seconds

and signal going to earth
D = 300,000 km/s * 0.005747 seconds
= 1724.1 km
Answer #2

2.0/348= 5.747*10^-3 seconds
How far does an electromagnetic wave travel in this time at 3*10^8m/s
Distance = 3*10^8*5.747*10^-3 =1724137.931m
The spaceship is 1,724km from earth